A high-rise building has forty floors.A cleaning company charge RM80 to clean the first floor and an extra RM40 for each floor above the preceding floor

a) find the cleaning for the twenty-fifth floor.
Answer #1

First floor charge = RM 80

Charges for the rest 24 floors = 24* 40 = 960

Total cleaning cost = 80+ 960 = 1040

Any doubt please comment, If it has helped you, please give a thumbs up.